Problem 13-6 Break-Even EBIT and Leverage [LO 1, 2] Silverton Co. is comparing two different capital structures. Plan I would result in 8,700 shares of stock and $323,000 in debt. Plan II would result in 12,000 shares of stock and $210,800 in debt. The interest rate on the debt is 10 percent.

a. Ignoring taxes, compare both of these plans to an all-equity plan assuming that EBIT will be $53,100. The all-equity plan would result in 18,200 shares of stock outstanding. Compute the EPS for each plan. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) EPS Plan I $ Plan II $ All-equity plan $

b. In part (a), what is the break-even level of EBIT for Plan I as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $ In part (a), what is the break-even level of EBIT for Plan II as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $

c. Ignoring taxes, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $

d. Assume the corporate tax rate is 30 percent. Compute the EPS for each plan. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) EPS Plan I $ Plan II $ All-equity plan $

What is the break-even level of EBIT for Plan I as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $

What is the break-even level of EBIT for Plan II as compared to that for an all-equity plan? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $

At what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) EBIT $
Answer #1

Answer a.

All-equity Plan:

Number of shares outstanding = 18,200

Plan I:

Number of shares outstanding = 8,700
Value of Debt = $323,000

Interest Expense = 10% * $323,000
Interest Expense = $32,300

Plan II:

Number of shares outstanding = 12,000
Value of Debt = $210,800

Interest Expense = 10% * $210,800
Interest Expense = $21,080

Plan I All-Equity Plan I 53,100 $ 32,300 $ 20,800 $ $ 20,800 $ |ЕВT Less: Interest |ЕВT Less: Taxes Net Income #of shares 53,

Answer b.

Let Breakeven EBIT be $x

All-equity Plan:

EPS = (EBIT - Interest Expense) / Number of shares outstanding
EPS = ($x - $0) / 18,200

Plan I:

EPS = (EBIT - Interest Expense) / Number of shares outstanding
EPS = ($x - $32,300) / 8,700

Plan II:

EPS = (EBIT - Interest Expense) / Number of shares outstanding
EPS = ($x - $21,080) / 12,000

Plan I and All-equity Plan:

EPS under Plan I and EPS under All-equity Plan
($x - $32,300) / 8,700 = ($x - $0) / 18,200
182 * $x - $5,878,600 = 87 * $x
$x = $61,880

Breakeven EBIT is $61,880

Plan II and All-equity Plan:

EPS under Plan II and EPS under All-equity Plan
($x - $21,080) / 12,000 = ($x - $0) / 18,200
182 * $x - $3,836,560 = 120 * $x
$x = $61,880

Breakeven EBIT is $61,880

Answer c.

Plan I and Plan II:

EPS under Plan I and EPS under Plan II
($x - $32,300) / 8,700 = ($x - $21,080) / 12,000
120 * $x - $3,876,000 = 87 * $x - $1,833,960
$x = $61,880

Breakeven EBIT is $61,880

Answer d-1.

All-Equity Plan I Plan II 53,100 $ $ 20,800 $ 6,240 $ 14,560 $ $ 53,100 $ ЕВT 53,100 Less: Interest $ 32,300 21,080 32,020 53

Answer d-2.

Let Breakeven EBIT be $x

All-equity Plan:

EPS = (EBIT - Interest Expense) * (1 - tax) / Number of shares outstanding
EPS = ($x - $0) * (1 - 0.30) / 18,200

Plan I:

EPS = (EBIT - Interest Expense) * (1 - tax) / Number of shares outstanding
EPS = ($x - $32,300) * (1 - 0.30) / 8,700

Plan II:

EPS = (EBIT - Interest Expense) * (1 - tax) / Number of shares outstanding
EPS = ($x - $21,080) * (1 - 0.30) / 12,000

Plan I and All-equity Plan:

EPS under Plan I and EPS under All-equity Plan
($x - $32,300) * 0.70 / 8,700 = ($x - $0) * 0.70 / 18,200
182 * $x - $5,878,600 = 87 * $x
$x = $61,880

Breakeven EBIT is $61,880

Plan II and All-equity Plan:

EPS under Plan II and EPS under All-equity Plan
($x - $21,080) * 0.70 / 12,000 = ($x - $0) * 0.70 / 18,200
182 * $x - $3,836,560 = 120 * $x
$x = $61,880

Breakeven EBIT is $61,880

Answer d-3.

Plan I and Plan II:

EPS under Plan I and EPS under Plan II
($x - $32,300) * 0.70 / 8,700 = ($x - $21,080) * 0.70 / 12,000
120 * $x - $3,876,000 = 87 * $x - $1,833,960
$x = $61,880

Breakeven EBIT is $61,880